OpenStudy (anonymous):

The difference of 32 and the quotient of a number m and 7 is 80Write an equation or inequality

OpenStudy (anonymous):

32-m/7=80

OpenStudy (anonymous):

32-m/7=80

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

or i think the more in depth way is |32-m/7|=80